# A Swift and Omnidirectional Formation Approach based on Hierarchical Reorganization

### v1

@@ Abstract
Current formations commonly rely on invariant hierarchical structures, such as predetermined leaders or enumerated formation shapes. These structures could be unidirectional and sluggish, constraining their adaptability and agility when encountering cluttered environments. To surmount these constraints, this work proposes an omnidirectional affine formation approach with hierarchical reorganizations. We first delineate the critical conditions requisite for facilitating hierarchical reorganizations within formations, which informs the development of the omnidirectional affine criterion. Central to our approach is the fluid leadership and authority redistribution, for which we develope a minimum time-driven leadership evaluation algorithm and a power transition control algorithm. These algorithms facilitate autonomous leader selection and ensure smooth power transitions, enabling the swarm to adapt hierarchically in alignment with the external environment. Furthermore, we deploy a power-centric topology switching mechanism tailored for the dynamic reorganization of in-team connections. Finally, simulations and experiments demonstrate the performance of the proposed method. The formation successfully performs several hierarchical reorganizations, with the longest reorganization taking only 0.047s. This swift adaptability allows five aerial robots to carry out complex tasks, including executing swerving movements and navigating through hoops at velocities up to 1.9m/s.
@@
Hierarchical Reorganization, Omnidirectional Affine Formation, Leader Selection, Swarm
@@ I Introduction
Aerial robots have garnered growing scholarly interest in recent years due to their vast potential for application in collaborative tasks, including exploration[1, 2], inspection[3], fire rescue[4]. Swarm formations, in particular, exhibit remarkable performance in self-organization, survivability, and collaborative task execution[5]. In complex environments, adopting variable formations during flight becomes crucial for maintaining optimal mobility and flexibility[6].
@@
The architectural frameworks facilitating formation reorganization can principally be categorized into democratic and autocratic. In democratic formations, the entire group is generally regarded as a unified entity, each planning their actions based on collective information and influencing the overall formation. For instance, in [7], the formation is conceptualized as a virtual rigid body, which allows for versatile transformations of various formation patterns by adjusting the relative orientations and positions of individual aerial robots with a virtual center. This approach facilitates the execution of complex tasks, such as navigating through cluttered buildings, dense forests, or disaster-stricken sites. However, this method relies on a virtual structure to govern the formation’s pattern, demanding real-time computation of transfer vectors for each individual within the formation. As the scale of aerial robots increases, the complexity of democratic reconfiguration correspondingly escalates, thereby diminishing the formation’s flexibility and robustness.
@@
Conversely, autocratic formation reorganization relies on a subset of agents within the formation. By hierarchically distributing the power among different individuals, certain members exert influence over the overall formation pattern, facilitating formation reorganization. For example, Ref. [8] introduces an affine formation method within a leader-follower framework. This method achieves comprehensive formation changes through the strategic positioning of the leader, enabling seamless continuous transitions, including translation, rotation, contraction, and deformation. However, the autocratic formations’ leaders remain unchanged, forming a unidirectional affine transformation. This configuration may pose challenges when the swarm needs to execute sharp turns or large-angle maneuvers, potentially limiting its flexibility.
@@
Given the abovementioned challenges, our approach draws inspiration from the swerving and hierarchical formation movements observed in starling clusters, as explored in [9, 10]. Starling flocks typically feature one or multiple leaders, with the formation’s internal framework exhibiting a structure between democratic and autocratic paradigms. Building on this, the cluster dynamically adjusts its movement patterns, leading to hierarchical reorganization within the formation. In light of this natural phenomenon, we introduce an omnidirectional affine formation, leveraging hierarchical reorganizations to enhance the flexibility of formations and in-team coordination.
Firstly, we elucidate the essential conditions for facilitating hierarchical reorganizations. Following this, we introduce a minimum time-driven leadership evaluation algorithm designed to streamline autonomous leader selection, thereby optimizing decision-making processes within the formation. With a focus on the seamless redistribution of leadership, a power transition control algorithm is proposed to ensure smooth authority redistribution. To enhance the formation’s internal synchrony, we deploy a power-centric topology switching mechanism tailored for the dynamic reorganization of in-team connections.
@@
    A novel omnidirectional affine formation strategy based on hierarchical reorganizations is proposed to efficiently improve the formation’s flexibility and coordination.
@@
    A power transition control algorithm is presented, which employs a designed power transfer function to achieve synchronized authority redistribution.
@@
The remainder of this paper is organized as follows. Related works are introduced in Section II. Preliminaries and problem formulation are given in Section III. The specific methodology is described in Section IV, which contains definitions for omnidirectional affine formation, a minimum time-driven leadership evaluation algorithm, a power transition control algorithm and a power-centric topology switch strategy. Verification is carried out in Section V, while Section VI concludes the paper.
@@ II Related Works
The related works are segmented into three sections. The first section explores algorithms dedicated to dynamic leader selection, reflecting on the existing criteria for fluid leadership. Subsequently, the second section delves into the methodologies governing hierarchical formation control, presenting a detailed overview of established approaches. The final section critically evaluates the role of topological switching, emphasizing its critical importance in hierarchical formation control.
II-A Dynamic Leader Selection
Researchers have made significant progress in leader selection to enhance various aspects of swarm formation performance [11, 12]. The focus is generally on selecting leaders based on specific task requirements. For instance, Ref.[13] formulates the leader selection problem as a fast mixed convex optimization problem, aiming to minimize convergence time by determining the appropriate number of leaders; Ref.[14] selects leaders based on criteria such as convergence speed and control energy. These studies underscore the importance of leader selection in optimizing swarm formation performance by considering factors like convergence time and task-specific requirements. However, these approaches are committed to static leader selection before initiating a single task. In practical situations, dynamic leader selection has been shown to improve the overall performance of a robot team compared to a static leader [15].
Dynamic leader selection involves changes in leader roles during the task due to various factors. On the one hand, this includes human or human base station influence over the robotic system. For instance, Ref.[16, 17, 18] improve communication quality with the base station by online leader selection in the swarm, and Ref.[19] proposes using human-robot trust as a dynamic criterion for leader selection, reducing task completion time and formation errors compared to non-leader switching strategies. However, these methods are constrained by the quality of human-machine communication and the accuracy of human judgment on the swarm’s current state, making them unsuitable for autonomous systems.
Considering the autonomy of the swarm, another aspect of dynamic leader selection is the influence of the external environment on the robotic swarm. For example, Ref.[20] suggests a method for reselecting leaders in case of leader failure in environments with dense obstacles. To address the leader trapping issue caused by different probabilities of encountering obstacles for different roles within the formation, Ref.[21] introduces an emotion-based model for leader selection. This model allows the team to autonomously reselect a leader when trapped and continue moving towards the goal. However, while this method can increase the probability of individual escape, it does not guarantee the successful escape of the entire swarm.
II-B Hierarchical Formation Control
@@
The hierarchical formation control can draw inspiration from collective movements observed in biological systems. For instance, the migration of bird flocks and mammal herds utilizes a leader-follower strategy to address the capability disparities among individuals [22, 23]. Numerous researchers have investigated formation control problems using leader-follower strategies [24, 25]. They have identified various formation structures, such as single-leader formation [26, 27, 28], multi-leader formation [29, 30], and virtual leader formation [31, 32]. Multi-leader structures, which allow for affine formation through adjusting leaders’ relative positions, have gained particular interest compared to single-leader and virtual-leader formation.
@@
To achieve affine transformations, Ref.[33] explores the sufficient and necessary conditions for realizing affine formations based on graph theory. Building on this, Ref.[8] proposes a control method for leader-follower formations based on affine transformation, demonstrating through simulations that the following error of followers converges exponentially. Furthering this approach, Ref.[34] introduces a hierarchical affine control algorithm to achieve formation control under conditions of non-global information, thereby enhancing the robustness of the control strategy. However, practical scenarios often involve communication delays among multiple agents, adversely affecting the effectiveness of the methods mentioned above.
Addressing the communication delays, several approaches have been proposed in recent research [35, 36]. One approach proposes using adaptive distributed observers to estimate the leader’s system matrix and state for each follower. By building a distributed controller, the approach achieves output consensus between leaders and followers [37]. Another approach focuses on the affine formation control of multi-agent systems with prescribed convergence times, which is accomplished by designing a distributed continuous control algorithm [38]. However, these control algorithms are all based on static and unidirectional structures, and research on formation control under hierarchical reorganizations is still in its infancy.
II-C Topology Switching for Hierarchical Formation
@@
Research has indicated that internal topology switching within a formation can enhance multiple capabilities of the formation. For example, Ref.[39] enhances the resilience of a swarm by designing internal topology transitions; Ref.[40] proposes an optimal rigidity graph-based topology optimization algorithm to reduce the communication complexity of formations, thereby extending the network’s lifespan; Ref.[41] designs smooth transitions in communication to ensure security during topological changes. The introduction of dynamic topology switching into formation control has added new variables [42, 43], making this area a subject of extensive research.
@@
Considering the control problems of formations under topology switching, Ref.[44] proposes the necessary and sufficient conditions for achieving time-varying formation tracking under topology transitions, and the system’s stability has been verified experimentally. Building on this, Ref.[45] presents a distributed control algorithm using neighboring positions and velocities during the topology transition process when weak connections exist in the communication topology among individuals in the swarm. Considering different formation structures, the problem of hierarchical formation control with topology switching has been further explored.
@@
To investigate the impact of topology switching on the control of hierarchical formation structures, Ref.[46] conducts a controllability analysis of hierarchical formations under topology switching. Based on this, Refs.[47, 48] study the control issues of leaders and followers during topology transitions. Furthermore, Refs.[33, 8] explore the relationship between hierarchical control issues in affine formation and topological structures. Considering the introduction of hierarchical reorganization strategies in these problems, Ref.[14] optimizes the formation of topological structures to reduce communication costs while improving convergence speed. However, no method has yet explored the impact of topology switching on formation control during hierarchical reorganizations.
@@ III-A Basic Graph and Formation Theory
Consider a group of $n$ mobile agents in ${\mathbf{R}}^{d}$ where $d>2$ and $n\geq d+1$. Let $\boldsymbol{p}_{i}\in\mathbf{R}^{d}$ be the position of agent $i$ and $\boldsymbol{p}=[\boldsymbol{p}_{1}^{T},...,\boldsymbol{p}_{n}^{T}]^{T}\in\mathbf{R}^{dn}$ be the configuration of all the agents. The interaction between the agents is modeled by a graph $\mathcal{G}(\mathcal{V},\mathcal{E})$ where $\mathcal{V}=\{1,2,...,n\}$ is the node set and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the edge set. For a digraph, the edge set consists of directed edges $(j,i)$, where node $j$ is the tail node and node $i$ is the head node. Node $j$ is called the in-neighbor of node $i$, while node $i$ is called the out-neighbor of node $j$ and $\mathcal{N}_{i}:=\{j|(j,i)\in\mathcal{E}\}$ denotes the in-neighbor set for agent $i$. In a digraph, $(j,i)\in\mathcal{E}\neq(i,j)\in\mathcal{E}$. An orientation of an undirected graph is the assignment of a direction to each undirected edge. An oriented graph is an undirected graph together with an orientation. The incidence matrix $\boldsymbol{D}\in\mathbf{R}^{m\times n}$ of an oriented graph is the $\{0,\pm 1\}$-matrix with rows indexed by edges and columns.
A formation, denoted as $(\mathcal{G},\boldsymbol{p})$, is the graph $\mathcal{G}$ with its vertex $i$ mapped to point $\boldsymbol{p}_{i}$. Without loss of generality, suppose the first $n_{l}$ agents are leaders and the rest $n_{f}=n-n_{l}$ agents are followers. Let $\mathcal{V}_{l}={1,...,n_{l}}$ and $\mathcal{V}_{f}=\mathcal{V}\setminus\mathcal{V}_{l}$ be the sets of leaders and followers, respectively. The positions of the leaders and followers are denoted as $\boldsymbol{p}_{l}=[\boldsymbol{p}_{1}^{T},...,\boldsymbol{p}_{n_{l}}^{T}]^{T}$ and $\boldsymbol{p_{f}}=[\boldsymbol{p}_{n_{l}+1}^{T},...,\boldsymbol{p}_{n}^{T}]^{T}$, respectively.
III-B Affine Formation Maneuver Control
For formation $(\mathcal{G},\boldsymbol{p})$, a stress is denoted as $\{\omega_{i}^{j}\}_{(i,j\in\mathcal{E})}$. According to [49], a stress is called an equilibrium stress if it satisfies

### v2

# leader_selection.bib

@@ Abstract
Conventional formation methods typically rely on fixed hierarchical structures, such as predetermined leaders or predefined formation shapes. These rigid hierarchies can render formations cumbersome and inflexible in complex environments, leading to potential failure if any leader loses connectivity. To address these limitations, this paper introduces a reconfigurable affine formation that enhances both flexibility and resilience through hierarchical reorganization. The paper first elucidates the critical role of hierarchical reorganization, conceptualizing this process as involving role reallocation and dynamic changes in topological structures. To further investigate the conditions necessary for hierarchical reorganization, a reconfigurable hierarchical formation is developed based on graph theory, with its feasibility rigorously demonstrated. In conjunction with role transitions, a power-centric topology switching mechanism grounded in formation consensus convergence is proposed, ensuring coordinated resilience within the formation. Finally, simulations and experiments validate the performance of the proposed method. The aerial formations successfully performed multiple hierarchical reorganizations in both three-dimensional and two-dimensional spaces. Even in the event of a single leader’s failure, the formation maintained stable flight through hierarchical reorganization. This rapid adaptability enables the robotic formations to execute complex tasks, including sharp turns and navigating through forests at speeds up to 1.9 m/s.
@@
Hierarchical Reorganization, Reconfigurable hierarchical formation, Leader Selection, Swarm
@@ I Introduction
Aerial robots have garnered growing scholarly interest in recent years due to their vast potential for application in collaborative tasks, including exploration[Hu2021, Zhou2022], inspection[Abdelkader2021], fire rescue[Merino2012]. Swarm formations, in particular, exhibit remarkable performance in self-organization, survivability, and collaborative task execution[Dong2015]. In complex environments, adopting variable formations during flight becomes crucial for maintaining optimal mobility and flexibility[Yu2022].
@@
The architectural frameworks facilitating formation reorganization can principally be categorized into democratic and autocratic. In democratic formations, the entire group is generally regarded as a unified entity, each planning their actions based on collective information and influencing the overall formation. For instance, in [Zhou2018], the formation is conceptualized as a virtual rigid body, which allows for versatile transformations of various formation patterns by adjusting the relative orientations and positions of individual aerial robots with a virtual center. This approach facilitates the execution of complex tasks, such as navigating through cluttered buildings, dense forests, or disaster-stricken sites. However, this method relies on a virtual structure to govern the formation’s pattern, demanding real-time computation of transfer vectors for each individual within the formation. As the scale of aerial robots increases, the complexity of democratic reconfiguration correspondingly escalates, thereby diminishing the formation’s flexibility and robustness.
@@
Conversely, autocratic formation reorganization relies on a subset of agents within the formation. By hierarchically distributing the power among different individuals, certain members exert influence over the overall formation pattern, facilitating formation reorganization. For example, Ref. [Zhao2018] introduces an affine formation method within a leader-follower framework. This method achieves comprehensive formation changes through the strategic positioning of the leader, enabling seamless continuous transitions, including translation, rotation, contraction, and deformation. However, the autocratic formations’ leaders remain unchanged, forming a unidirectional affine transformation. This configuration may pose challenges when the swarm needs to execute sharp turns or large-angle maneuvers, potentially limiting its flexibility. Notably, the entire formation with fixed hierarchy is at risk of collapse if any single leader becomes inoperative.
@@
Given the abovementioned challenges, our approach draws inspiration from the swerving and hierarchical formation movements observed in starling clusters, as explored in [Mateo2017, Attanasi2014]. Starling flocks typically feature one or multiple leaders, with the formation’s internal framework exhibiting a structure between democratic and autocratic paradigms. Building on this, the cluster dynamically adjusts its movement patterns, leading to hierarchical reorganization within the formation. In light of this natural phenomenon, we introduce an reconfigurable hierarchical formation(RHF), leveraging hierarchical reorganizations to enhance the flexibility and resilience of formation.
First, we elucidate the definition of hierarchical reorganization for a formation. Building on this, the paper defines a reconfigurable hierarchical formation based on affine formation theory and analyzes the configuration requirements necessary for achieving hierarchical reorganization, providing rigorous proof. Subsequently, to enhance the internal synchronization of dynamic formations, a power-centric topology switching mechanism is proposed based on formation consensus convergence. This mechanism adapts to the dynamic reorganization of intra-formation connections, ensuring coordinated resilience of the formation.
@@
    A novel RHF strategy is proposed to efficiently improve the formation’s flexibility and resilience.
@@
    The necessary conditions and theory analysis for hierarchical reorganization are proposed, with their feasibility and necessity rigorously demonstrated.
@@
  4. 4.
    Simulation and Experiments are carried out to verify the performance of the RHF strategy.
The remainder of this paper is organized as follows. Related works are introduced in Section II. Preliminaries and problem formulation are given in Section III. The specific methodology is described in Section IV, which contains specific analysis of RHF and a power-centric topology switch strategy. Verification is carried out in Section V, while Section VI concludes the paper.
@@ II Related Works
The related works are segmented into two sections. The first section explores hierarchical affine formation, including the development of affine formations and the current state of dynamic formation development. The second section critically evaluates the role of topological switching, emphasizing its critical importance in hierarchical formation.
II-A Hierarchical Affine Formation
@@
Hierarchical formation control, with its advantage of addressing capability disparities among individuals, has become one of the primary methods in formation control[Chen2015]. Numerous researchers have investigated formation control problems using leader-follower strategies [Cao2013, Oh2015]. They have identified various formation structures, such as single-leader formation[Kang2014, Babazadeh2018, Dai2022], multi-leader formation[Fang2022, Nomura2021], and virtual leader formation[Porfiri2007, Droge2015]. Multi-leader structures, which allow for affine formation through adjusting leaders’ relative positions, have gained particular interest compared to single-leader and virtual-leader formation.
@@
To achieve affine transformations, Ref.[Lin2016] explores the sufficient and necessary conditions for realizing affine formations based on graph theory. Building on this, Ref.[Zhao2018] proposes a control method for leader-follower formations based on affine transformation, demonstrating through simulations that the following error of followers converges exponentially. Furthering this approach, Ref.[Li2021] introduces a hierarchical affine control algorithm to achieve formation control under conditions of non-global information, thereby enhancing the robustness of the control strategy. However, these control algorithms are all based on static and predetermined structures, which may be unresponsive and cumbersome, making it difficult for the formation to perform large maneuvers while maintaining formation in complex environments.
Based on this, researchers have found that in practical situations, dynamic leader selection has been shown to improve the overall performance of a robot team compared to a static leader [Franchi2011]. Current dynamic leader selection involves changes in leader roles during the task due to various factors. On the one hand, this includes human or human base station influence over the robotic system. For instance, Ref.[Misra2021, AbdelMalek2022, Ganesan2020] improve communication quality with the base station by online leader selection in the swarm, and Ref.[Saeidi2017] proposes using human-robot trust as a dynamic criterion for leader selection, reducing task completion time and formation errors compared to non-leader switching strategies. However, these methods are constrained by the quality of human-machine communication and the accuracy of human judgment on the swarm’s current state, making them unsuitable for autonomous systems.
Considering the autonomy of the swarm, another aspect of dynamic leader selection is the influence of the external environment on the robotic swarm. For example, Ref.[Tavares2020] suggests a method for reselecting leaders in case of leader failure in environments with dense obstacles. To address the leader trapping issue caused by different probabilities of encountering obstacles for different roles within the formation, Ref.[Li2017] introduces an emotion-based model for leader selection. This model allows the team to autonomously reselect a leader when trapped and continue moving towards the goal. However, while this method can increase the probability of individual escape, it does not guarantee the successful escape of the entire swarm.
II-B Topology Switching for Hierarchical Formation
@@
Research has indicated that internal topology switching within a formation can enhance multiple capabilities of the formation. For example, Ref.[Yu2022a] enhances the resilience of a swarm by designing internal topology transitions; Ref.[Zhao2023] proposes an optimal rigidity graph-based topology optimization algorithm to reduce the communication complexity of formations, thereby extending the network’s lifespan; Ref.[Yu2020a] designs smooth transitions in communication to ensure security during topological changes. The introduction of dynamic topology switching into formation control has added new variables [Mesbahi2010, OlfatiSaber2007], making this area a subject of extensive research.
@@
Considering the control problems of formations under topology switching, Ref.[Dong2017] proposes the necessary and sufficient conditions for achieving time-varying formation tracking under topology transitions, and the system’s stability has been verified experimentally. Building on this, Ref.[Zou2018] presents a distributed control algorithm using neighboring positions and velocities during the topology transition process when weak connections exist in the communication topology among individuals in the swarm. Considering different formation structures, the problem of hierarchical formation control with topology switching has been further explored.
@@
To investigate the impact of topology switching on the control of hierarchical formation structures, Ref.[Liu2008] conducts a controllability analysis of hierarchical formations under topology switching. Based on this, Refs.[Wang2012, Soni2021] study the control issues of leaders and followers during topology transitions. Furthermore, Refs.[Lin2016, Zhao2018] explore the relationship between hierarchical control issues in affine formation and topological structures. Considering the introduction of hierarchical reorganization strategies in these problems, Ref.[Xiao2022] optimizes the formation of topological structures to reduce communication costs while improving convergence speed. However, no method has yet explored the impact of topology switching on formation control during hierarchical reorganizations.
@@ III-A Basic Graph and Formation Theory
Consider a group of $n$ mobile agents in ${\mathbb{R}}^{d}$ where $d>2$ and $n\geq d+1$. Let $\boldsymbol{p}_{i}\in\mathbb{R}^{d}$ be the position of agent $i$ and $\boldsymbol{p}=[\boldsymbol{p}_{1}^{T},...,\boldsymbol{p}_{n}^{T}]^{T}\in\mathbb{R}^{dn}$ be the configuration of all the agents. The interaction between the agents is modeled by a graph $\mathcal{G}(\mathcal{V},\mathcal{E})$ where $\mathcal{V}=\{1,2,...,n\}$ is the node set and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the edge set. For a digraph, the edge set consists of directed edges $(j,i)$, where node $j$ is the tail node and node $i$ is the head node. Node $j$ is called the in-neighbor of node $i$, while node $i$ is called the out-neighbor of node $j$ and $\mathcal{N}_{i}:=\{j|(j,i)\in\mathcal{E}\}$ denotes the in-neighbor set for agent $i$. In a digraph, $(j,i)\in\mathcal{E}\neq(i,j)\in\mathcal{E}$.
A formation, denoted as $(\mathcal{G},\boldsymbol{r})$, is the graph $\mathcal{G}$ with its vertex $i$ mapped to point $\boldsymbol{r}_{i}$, where $\boldsymbol{r}$ represents the configuration of $\mathcal{G}$. Without loss of generality, suppose the first $n_{l}$ agents are leaders and the rest $n_{f}=n-n_{l}$ agents are followers. Let $\mathcal{V}_{l}={1,...,n_{l}}$ and $\mathcal{V}_{f}=\mathcal{V}\setminus\mathcal{V}_{l}$ be the sets of leaders and followers, respectively. The current positions of the leaders and followers are denoted as $\boldsymbol{p}_{l}=[\boldsymbol{p}_{1}^{T},...,\boldsymbol{p}_{n_{l}}^{T}]^{T}$ and $\boldsymbol{p_{f}}=[\boldsymbol{p}_{n_{l}+1}^{T},...,\boldsymbol{p}_{n}^{T}]^{T}$, respectively.
III-B Affine Image and Affine Localizability
For configuration $\boldsymbol{r}$, if there are real matrixs $\mathbb{A}$ and $\mathbb{B}$ with appropriate dimensions, $\mathcal{A}(\boldsymbol{r})=\{\boldsymbol{p}\in\mathbb{R}^{dn}:=(\boldsymbol{I}_{n}\otimes\mathbb{A})\boldsymbol{r}+\mathbb{1}\otimes\boldsymbol{B}\}$, $\mathcal{A}(\boldsymbol{r})$ is defined as an Affine Image of $\boldsymbol{p}$. Given that the formation configuration $\boldsymbol{r}$, then at any time during the affine formation movement, the formation $\boldsymbol{p}$ must satisfy that $\boldsymbol{p}\in\mathcal{A}(\boldsymbol{r})$.
In a hierarchical formation, it is necessary to infer the current position of the followers based on the position of the leader. According to [Zhao2018], affine localizability can be defined as follows: For any $\boldsymbol{p}=[\boldsymbol{p}_{l},\boldsymbol{p}_{f}]^{T}\in\mathcal{A}(r)$, if $\boldsymbol{p}_{f}$ can be uniquely determined by $\boldsymbol{p}_{l}$, then the formation is affinely localizable.
The conditions for affine localizability can be divided into position and stress conditions. Given a set of points $\boldsymbol{p}$, let $\bar{\boldsymbol{p}}\in\mathbb{R}^{n\times(d+1)}=[\boldsymbol{p},\mathbb{1}_{n}]$. If and only if $n\geq d+1$ and $rank(\bar{\boldsymbol{p}})=d+1$, $\boldsymbol{p}$ affinely span $\mathbb{R}^{d}$. For the formation $(\mathcal{G},\boldsymbol{r})$, assume that $\boldsymbol{r}$ affinely span $\mathbb{R}^{d}$. The position condition for affine localizability is as follows:
Lemma III.1.
The formation $(\mathcal{G},\boldsymbol{r})$ if affinely localizability if and only if $\boldsymbol{r}$ affinely span $\mathbb{R}^{d}$.
III-C Stress Matrices and Affine Maneuver
For formation $(\mathcal{G},\boldsymbol{r})$, a stress is denoted as $\{\omega_{i}^{j}\}_{(i,j\in\mathcal{E})}$. According to Ref. [Connelly2005], a stress is called an equilibrium stress if it satisfies